OpenStudy (anonymous):

can anyone help me set up the equation of: Cashews are priced at $5.75 per pound. Peanuts are priced at $3.25 per pound. Mixed together to total 80 pounds for $4.00 per pound. How many pounds of each do I need?

OpenStudy (anonymous):

x+y=80 3.25x+5.75y=4.00 Pretty sure that is right just use the elimination or substitution methode to figure it out

OpenStudy (anonymous):

Let me know if it works out for you i am wondering myself now

OpenStudy (anonymous):

my class is not to substitution or elimination. i was thinking: 3.25x+5.75(x-80)=320?

OpenStudy (anonymous):

i don't know how you would solve that if you are not using one of those methodes. to me it looks like it should be a set of equations in multi varible form. let me think of this for a sec.

OpenStudy (anonymous):

I have tutoring before class tomorrow, I was just trying to figure it out before then. No worries. thank you.

OpenStudy (amistre64):

its helpful to use variables that represent the actual stuff :) c for Cashew and p for peanuts helps you to keep track of them better ;) other than that you did a great job with it

OpenStudy (amistre64):

c+p=80 p(3.25)+c(5.75)=(80)4.00

OpenStudy (amistre64):

solve for one variable in the first equation: c = 80-p and substitute it into the second.... p(3.25)+(80-p)(5.75)=(80)4.00 do the math, and solve for p..

OpenStudy (amistre64):

once you know a solid value for "p" use it in the c = 80-p to find c.